OpenStudy (anonymous):

An initial deposit of $20,000 grows at an annual rate of 4% for 14 years. Compare the final balances resulting from annual compounding and continuous compounding. (Give your answers correct to the nearest cent.) Explain Answers please

OpenStudy (phi):

Here is the start of some videos that might help you, if you need more info on how to do this stuff http://www.khanacademy.org/finance-economics/core-finance/v/introduction-to-interest

OpenStudy (phi):

You need to know some formulas, and how to use them. For Compound interest, use \[ FutureValue= PresentValue \cdot(1+\frac{i}{n})^{nt} \] where i is the interest rate (as a decimal, e.g. 4% = 0.04) n is the number of times compounded per year. With your problem, n=1 (compounded annually means compounded once a year) t is the number of years. For your problem, t= 14 years Your Present Value = 20,000. So if we plug in the numbers that we know, i= 0.04, n=1, t=14, you must solve \[ FV= 20000\cdot (1.04)^{14} \] You will need a calculator to do this.

OpenStudy (phi):

the formula for continuous compounding is: \[ FV= PV \cdot e^{i\cdot t} \] where i is the interest rate= 0.04, and t= the years= 14 years

OpenStudy (anonymous):

\[ 20000. e^{0.04 \times 14}-20000 (1+0.04)^{14}=379.921 \]